# DynOPETs: A Versatile Benchmark for Dynamic Object Pose Estimation and Tracking in Moving Camera Scenarios

### v1

@@ IV EVALUATION
IV-A Overview
@@
As discussed in  section II-A1, IOPE methods are not our primary focus, instead, we adopt the COPE and UOPE methods for benchmarking, which are capable of generalizing across intra-class variations and unseen objects without the need for individual retraining for each object. These methods can directly leverage open-source pre-trained weights. This approach serves two purposes: first, it allows us to verify the validity and accuracy of our data; second, it helps evaluate the generalization capability of the open-source algorithm on our dataset. Furthermore, we fine-tuned several variants of the COPE method on the COPE-119 training set and evaluated their performance on the test set to validate the effectiveness of our ground truth labels for network training. Additionally, we assessed various object-pose-tracking methods to further broaden our analysis.
IV-B Evaluation Metrics
@@
We follow [8] and adopt the evaluation protocol commonly used for COPE methods, which includes the mean precision of 3D intersection over union (IoU) to jointly assess rotation, translation, and size prediction. 3D IoU measures the overlap between the 3D bounding box of an object derived from the predicted pose and the ground truth bounding box, providing a comprehensive metric. For rotation and translation errors, we typically report the accuracy within a specified threshold. For instance, RotmTransn represents the percentage of predicted poses with a rotation error less than $m$ degrees and a translation error less than $n$ centimeters.
@@
We evaluated the Average Recall (AR) [50] for UOPE methods using three error functions: Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD). In addition, we report the frame rate (FPS) for all UOPE methods.
@@
The evaluation of OPT methods performance relies on two complementary metrics: We evaluate object pose tracking performance using Absolute Trajectory Error (ATE) for global accuracy and Relative Pose Error (RPE) [51] for local accuracy, reporting average results across all 175 sequences. We also report the AUC of ADD and ADD-S [11], as well as the average recall of ADD(S) within 10% of the object diameter (termed ADD(S)-0.1d) [52].
IV-C Category-level Pose Estimation
IV-C1 COPE-119 Baseline Evaluation
@@
For COPE task, we evaluate 8 methods:  [8],  [49], and  [32] predict canonical representations directly, while  [39],  [35],  [37],  [40], and  [41] predict object pose and category. Except for NOCS [8], which performs joint object detection and pose estimation, all methods utilize identical SAM2 [45] segmentation results to ensure a fair benchmark. Notably, DiffusionNOCS [32] achieves a relatively lower IoU score but exhibits a very small pose error. A possible reason is that it has never been trained on real-world data, leading to instability in object scale prediction.
IV-C2 COPE-119 Fine-Tuning Evaluation
@@
The purpose of this experiment is to demonstrate the accuracy of pose annotations, based on the assumption that the metric improves only when the ground truth pose annotations in both the training and test sets are correct. We fine-tuned AGPose [41] and SecondPose [40] on our 84-sequence training set for 2 epochs and evaluated them on our 35-sequence test set. As shown in  table III, the fine-tuned results significantly outperform their original test set performance by a large margin in merely 2 epochs of fine-tuning, highlighting the accuracy of pose annotations in our dataset.
@@
Method Input $\uparrow$ VSD $\uparrow$ MSSD $\uparrow$ MSPD $\uparrow$ AR $\uparrow$ FPS $\downarrow$
FoundPose* [30] RGB 50.12 52.02 79.25 60.46 0.50
MegaPose [28] RGB 64.56 68.45 83.64 72.23 2.25
GigaPose [29] RGB 64.84 61.32 83.20 69.79 3.37
MegaPose [28] RGB-D 76.34 80.48 84.64 80.49 2.61
SAM6D [25] RGB-D 79.67 77.95 81.65 79.75 3.75
FoundationPose [1] RGB-D 92.74 90.09 94.40 92.41 0.98
IV-D Unseen Object Pose Estimation
We evaluate 5 UOPE methods in table table IV, grouped by their input modalities into RGB-based and RGB-D-based approaches. To ensure fairness, each sequence is paired with a specific CAD model and uniform segmentations produced by SAM2 [45]. Moreover, methods that rely on CNOS [53] for detection or segmentation also utilize SAM2. Except for SAM6D [25], which experiences performance bottlenecks due to its intrinsic segmentation component and fixed threshold settings that fail to generalize in our dataset’s cluttered backgrounds, we apply dynamic threshold adjustments to maintain accurate pose estimation. MegaPose [28], available in RGB-only and RGB-D configurations, leverages extensive pre-processing to extract rich template information, demonstrating a significant advantage, especially on single-object sequences with ground-truth detection. Notably, FoundPose [30] does not open-source its core refinement code, limiting its evaluation of the results, so we only report its coarse results.
IV-E Object Pose Tracking
For object pose tracking methods, we benchmark  [54],  [42],  [4],  [43], and  [30]. We mainly employed the original algorithms, but with DROID-SLAM [7], we integrated SAM2-derived masks [45] to zero out weights of non-object regions directly within the DBA layer to derive the camera-to-object poses, effectively adapting it for object pose tracking. We report FoundationPose [1] results under the model-based mode.
@@ IV-E Object Pose Tracking
MaskFusion [54] 0.320 0.060 0.033
DROID-SLAM [4] 0.173 0.063 0.040
BundleTrack [42] 0.068 0.037 0.021
BundleSDF [43] 0.094 0.044 0.026
FoundationPose [1] 0.037 0.018 0.010
@@
MaskFusion [54] 12.23 34.47 28.75 64.65
Droid-SLAM [4] 34.27 57.20 62.20 86.41
BundleTrack [42] 76.39 89.36 94.30 99.54
BundleSDF [43] 70.64 89.41 92.30 99.98
FoundationPose [1] 89.80 96.40 92.42 99.29
By considering both ATE and RPE, we obtain a comprehensive understanding of the pose estimation performance in table table V, encompassing both global consistency and local accuracy. Besides, we benchmark each method’s performance in table table VI across all 175 sequences (COPE-119 & UOPE-56) using the AUC of ADD(S) and ADD(S)-0.1d. Due to failure cases in some sequences, FoundationPose [1] underperforms—especially in ADD(S)-0.1d compared to BundleTrack [42] and BundleSDF [43]. Similarly, MaskFusion [54] shows poor performance, consistent with the trajectory error metrics.

### v2

@@ IV EVALUATION
IV-A Comparsions with Pose from Motion Capture System
Before conducting benchmarks, we need to verify the accuracy of the pose labels obtained from the proposed annotation algorithm. For this purpose, we collected an additional 10 sequences of data, where the objects in these sequences were equipped with motion capture system markers, allowing us to obtain ground truth with millimeter-level accuracy. We use the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) [sturm2012benchmark] to evaluate the pose annotation compared to the ground truth object pose. The results show that the average rotation error between our poses and those obtained from the motion capture system is less than one degree, and the average translation error is less than 1 cm, which is sufficiently accurate. Due to space limitations, please refer to the PDF supplementary material on the project page for details of this experiment.
IV-B Overview
@@
As discussed in  section II-A1, IOPE methods are not our primary focus, instead, we adopt the COPE and UOPE methods for benchmarking, which are capable of generalizing across intra-class variations and unseen objects without the need for individual retraining for each object. These methods can directly leverage open-source pre-trained weights. This approach serves two purposes: first, it allows us to verify the validity and accuracy of our data; second, it helps evaluate the generalization capability of the open-source algorithm on our dataset. Furthermore, we fine-tuned several variants of the COPE method on the COPE-119 training set and evaluated their performance on the test set to validate the effectiveness of our ground truth labels for network training. Additionally, we assessed various object-pose-tracking methods to further broaden our analysis. Finally, it is important to note that due to space limitations, detailed experimental analysis and in-depth discussion can be found in the PDF supplementary material on the project page.
IV-C Evaluation Metrics
@@
We follow [wang2019normalized] and adopt the evaluation protocol commonly used for COPE methods, which includes the mean precision of 3D intersection over union (IoU) to jointly assess rotation, translation, and size prediction. 3D IoU measures the overlap between the 3D bounding box of an object derived from the predicted pose and the ground truth bounding box, providing a comprehensive metric. For rotation and translation errors, we typically report the accuracy within a specified threshold. For instance, RotmTransn represents the percentage of predicted poses with a rotation error less than $m$ degrees and a translation error less than $n$ centimeters.
@@
We evaluated the Average Recall (AR) [hodan2023bop] for UOPE methods using three error functions: Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD). In addition, we report the frame rate (FPS) for all UOPE methods.
@@
The evaluation of OPT methods performance relies on two complementary metrics: We evaluate object pose tracking performance using ATE for global accuracy and RPE [sturm2012benchmark] for local accuracy, reporting average results across all 175 sequences. We also report the AUC of ADD and ADD-S [xiang2018posecnn], as well as the average recall of ADD(S) within 10% of the object diameter (termed ADD(S)-0.1d) [he2022oneposeplusplus].
IV-D Category-level Pose Estimation
IV-D1 COPE-119 Baseline Evaluation
@@
As shown in table II, for COPE task, we evaluate 8 methods:  [wang2019normalized],  [yang2024mv], and  [ikeda2024diffusionnocs] predict canonical representations directly, while [li2023generative]  [lin2023vi],  [liu2023net],  [zhang2023generative],  [chen2024secondpose], and  [lin2024instance] predict object pose and category. Except for NOCS [wang2019normalized], which performs joint object detection and pose estimation, all methods utilize identical SAM2 [ravi2024sam] segmentation results to ensure a fair benchmark. Notably, DiffusionNOCS [ikeda2024diffusionnocs] achieves a relatively lower IoU score but exhibits a very small pose error. A possible reason is that it has never been trained on real-world data, leading to instability in object scale prediction.
IV-D2 COPE-119 Fine-Tuning Evaluation
@@
The purpose of this experiment is to demonstrate the accuracy of pose annotations, based on the assumption that the metric improves only when the ground truth pose annotations in both the training and test sets are correct. We fine-tuned AGPose [lin2024instance] and SecondPose [chen2024secondpose] on our 84-sequence training set for 2 epochs and evaluated them on our 35-sequence test set. As shown in  table III, the fine-tuned results significantly outperform their original test set performance by a large margin in merely 2 epochs of fine-tuning, highlighting the accuracy of pose annotations in our dataset.
@@
Method Input VSD $\uparrow$ MSSD $\uparrow$ MSPD $\uparrow$ AR $\uparrow$ FPS $\downarrow$
FoundPose* [ornek2024foundpose] RGB 50.12 52.02 79.25 60.46 0.50
MegaPose [labbe2022megapose] RGB 64.56 68.45 83.64 72.23 2.25
GigaPose [nguyen2024gigapose] RGB 64.84 61.32 83.20 69.79 3.37
MegaPose [labbe2022megapose] RGB-D 76.34 80.48 84.64 80.49 2.61
SAM6D [lin2024sam] RGB-D 79.67 77.95 81.65 79.75 3.75
FoundationPose [wen2024foundationpose] RGB-D 92.74 90.09 94.40 92.41 0.98
IV-E Unseen Object Pose Estimation
We evaluate 5 UOPE methods in table IV, grouped by their input modalities into RGB-based and RGB-D-based approaches. To ensure fairness, each sequence is paired with a specific CAD model and uniform segmentations produced by SAM2 [ravi2024sam]. Moreover, methods that rely on CNOS [nguyen2023cnos] for detection or segmentation also utilize SAM2. Except for SAM6D [lin2024sam], which experiences performance bottlenecks due to its intrinsic segmentation component and fixed threshold settings that fail to generalize in our dataset’s cluttered backgrounds, we apply dynamic threshold adjustments to maintain accurate pose estimation. MegaPose [labbe2022megapose], available in RGB-only and RGB-D configurations, leverages extensive pre-processing to extract rich template information, demonstrating a significant advantage, especially on single-object sequences with ground-truth detection. Notably, FoundPose [ornek2024foundpose] does not open-source its core refinement code, limiting its evaluation of the results, so we only report its coarse results.
IV-F Object Pose Tracking
For object pose tracking methods, we benchmark  [runz2018maskfusion],  [wen2021bundletrack],  [teed2021droid],  [wen2023bundlesdf], and  [ornek2024foundpose]. We mainly employed the original algorithms, but with DROID-SLAM [teed2021droid], we integrated SAM2-derived masks [ravi2024sam] to zero out weights of non-object regions directly within the DBA layer to derive the camera-to-object poses, effectively adapting it for object pose tracking. We report FoundationPose [wen2024foundationpose] results under the model-based mode.
@@ IV-F Object Pose Tracking
MaskFusion [runz2018maskfusion] 0.320 0.060 0.033
DROID-SLAM [teed2021droid] 0.173 0.063 0.040
BundleTrack [wen2021bundletrack] 0.068 0.037 0.021
BundleSDF [wen2023bundlesdf] 0.094 0.044 0.026
FoundationPose [wen2024foundationpose] 0.037 0.018 0.010
@@
MaskFusion [runz2018maskfusion] 12.23 34.47 28.75 64.65
DROID-SLAM [teed2021droid] 34.27 57.20 62.20 86.41
BundleTrack [wen2021bundletrack] 76.39 89.36 94.30 99.54
BundleSDF [wen2023bundlesdf] 70.64 89.41 92.30 99.98
FoundationPose [wen2024foundationpose] 89.80 96.40 92.42 99.29
By considering both ATE and RPE, we obtain a comprehensive understanding of the pose estimation performance in table table V, encompassing both global consistency and local accuracy. Besides, we benchmark each method’s performance in table table VI across all 175 sequences (COPE-119 & UOPE-56) using the AUC of ADD(S) and ADD(S)-0.1d. Due to failure cases in some sequences, FoundationPose [wen2024foundationpose] underperforms—especially in ADD(S)-0.1d compared to BundleTrack [wen2021bundletrack] and BundleSDF [wen2023bundlesdf]. Similarly, MaskFusion [runz2018maskfusion] shows poor performance, consistent with the trajectory error metrics.
@@ V CONCLUSION
[…]
[heading=bibintoc,title=References]
{refsection}
VI Supplementary
VI-A Experiments with Motion Capture System
VI-A1 Experiment Setup
In addition to the original 175 sequences in our dataset, we collected 10 extra validation sequences shown in fig. 4 with precise object pose ground truth using a motion capture system. We made every effort to ensure these 10 additional sequences closely mirror the experimental conditions of the original 175 sequences. These 10 sequences feature 10 different objects: 6 objects come from 6 categories in COPE-119, and 4 objects are from UOPE-56. To provide ground truth poses using the motion capture system, each object was affixed with at least three markers, forming a rigid body frame within the motion capture system coordinate system. We refer to the rigid body frames formed by markers attached to the object and the camera as the object body frame and camera body frame, respectively. The position and orientation of object body in the motion capture system is denoted as $\textbf{T}_{OB,M}$. It is important to note that the orientation and position of this object body frame differ from those of the object’s own coordinate system; there is a fixed but unknown transformation $\textbf{T}_{O,OB}$ between them. We determined this unknown transformation based on the geometric relationship
$\textbf{T}_{O,OB}=(\textbf{T}_{OB,M})^{-1}\textbf{T}_{CB,M}\textbf{T}_{C,CB}\textbf{T}_{O,C}$ (10)
, where $\textbf{T}_{C,CB}$ is the handeye calibration result representing the transformation from camera to camera body frame, $\textbf{T}_{CB,M},\textbf{T}_{OB,M}$ are motion capture reading of camera body and object body position and orentation in motion capture system world coordinate, and $\textbf{T}^{O}_{C}$ is the object pose which can be estimated by algorithms like FoundationPose [wen2024foundationpose]. Note that, to ensure accuracy, in our setup, $\textbf{T}_{O,C}$ is estimated by  [wen2024foundationpose] first then refined by manual alignment of the object point cloud and its CAD model. Finally, the ground truth object pose is given by
$\textbf{T}_{O,C}=(\textbf{T}_{C,CB})^{-1}(\textbf{T}_{CB,M})^{-1}\textbf{T}_{OB,M}\textbf{T}_{O,OB}$ (11)
. Please note that after the markers are attached to these 10 objects, we reconstructed object CAD models using an RGB-D scanner. However, because the markers themselves reflect infrared light, they interfered with the RGB-D scanner’s depth readings, leading to imperfections in certain areas of the object mesh. Despite this, the overall quality remains usable.
Figure 4: The 10 objects in Mocap experiments we used.
VI-A2 Implementation Details
We benchmark several key outputs against the ground truth object pose: the raw, unprocessed FoundationPose [wen2024foundationpose] (referred to as Raw Abs-Pose), the refined and smoothed  [wen2024foundationpose] by global Kalman Filter (Refined Abs-Pose), as well as the outputs of our relative pose estimator (Rel-Pose) and the pose graph optimization results (PGO-Pose). In pose graph optimization, the information matrix of the absolute pose edge is set heuristically. First, for each frame, we reproject the object bounding box and axis onto the 2D image based on the Refined Abs-Pose estimation. Then, we stitch these images into a video. By observing whether the projected bounding box in the video exhibits jitter, we can assess the quality of the absolute pose estimation. Typically, 2D projection jitter indicates an inconsistency between the object poses in consecutive frames. This means that the object poses across the few frames exhibiting this jitter will inevitably vary in quality—some may be good, some bad, or all of them might be poor. Consequently, these poses are considered unreliable.
In general, the quality of absolute pose estimation is better than that of relative pose estimation. Therefore, we adopt a conservative parameter strategy, setting the diagonal elements of the absolute pose information matrix to 1e5 to reduce the negative impact of the relative pose edge.
For those visually unreliable absolute edges, the information matrix elements are set to 1e2-1e3. Note that the elements of the relative pose edge information matrix are generally around 1e2. This allows the relative pose to effectively compensate for poor absolute pose estimations. Finnally, for our comparative analysis, we utilize ATE (Absolute Trajectory Error) and RPE (Relative Pose Error). The unit of translation error is millimeters and the unit of rotation error is degrees.
TABLE VII: ATE (mm) results.
Sequence ID Rel-Pose Raw Abs-Pose Refined Abs-Pose PGO Pose
ATE. ATE. ATE. ATE. ATE. ATE. ATE. ATE. ATE. ATE. ATE. ATE.
max mean median max mean median max mean median max mean median
Seq0 484.70 159.37 146.14 24.17 6.15 5.60 16.85 5.98 5.56 16.85 5.99 5.56
Seq1 364.25 176.07 166.79 17.01 4.90 4.84 17.01 4.89 4.86 17.01 4.88 4.86
Seq2 534.79 274.89 280.95 17.16 5.15 4.92 17.16 5.06 4.78 17.16 5.07 4.78
Seq3 228.37 85.35 83.48 16.79 6.35 5.37 16.79 6.35 5.38 16.79 6.35 5.38
Seq4 357.11 165.51 160.60 16.87 5.91 5.84 16.88 5.92 5.84 16.88 5.93 5.84
Seq5 257.58 111.89 90.20 7.59 3.14 3.00 7.59 3.15 3.01 7.59 3.15 3.01
Seq6 209.48 105.43 99.38 16.35 5.70 5.38 16.34 5.71 5.39 16.35 5.72 5.41
Seq7 250.53 117.48 121.19 19.91 4.86 3.93 18.48 4.71 3.87 18.50 4.71 3.87
Seq8 292.44 92.27 83.14 12.74 3.72 3.47 12.73 3.71 3.46 12.73 3.71 3.46
Seq9 296.40 153.10 165.70 13.95 3.90 3.33 13.95 3.85 3.25 13.95 3.85 3.25
Avg. 327.37 144.62 139.76 16.65 4.98 4.53 15.38 4.93 4.54 15.38 4.94 4.54
TABLE VIII: RPE results for the rot. (degs).
Sequence ID Rel-Pose Raw Abs-Pose Refined Abs-Pose PGO Pose
RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE.
rot.max rot.mean rot.median rot.max rot.mean rot.median rot.max rot.mean rot.median rot.max rot.mean rot.median
Seq0 5.31 1.15 0.80 5.54 1.12 0.89 3.24 0.21 0.14 3.21 0.21 0.14
Seq1 7.54 1.05 0.78 9.80 0.95 0.75 5.36 0.15 0.10 3.69 0.14 0.10
Seq2 139.58 1.90 0.86 2.95 0.97 0.87 0.45 0.13 0.12 0.71 0.13 0.12
Seq3 8.71 0.96 0.80 5.64 1.31 1.11 1.60 0.43 0.33 2.88 0.43 0.33
Seq4 3.59 0.63 0.52 3.00 0.76 0.63 2.86 0.36 0.27 2.86 0.36 0.27
Seq5 3.21 0.77 0.64 3.29 0.83 0.72 0.26 0.10 0.09 0.26 0.10 0.09
Seq6 2.91 0.72 0.59 5.69 0.81 0.58 2.37 0.14 0.09 1.44 0.14 0.09
Seq7 3.86 0.78 0.64 4.94 0.80 0.61 0.57 0.13 0.11 0.55 0.13 0.11
Seq8 3.37 0.74 0.68 2.50 0.73 0.62 0.52 0.11 0.09 0.49 0.11 0.09
Seq9 5.44 0.85 0.58 3.97 0.79 0.66 2.78 0.16 0.13 1.77 0.16 0.13
Avg. 18.05 0.95 0.69 4.93 0.91 0.74 2.00 0.19 0.15 1.79 0.19 0.15
TABLE IX: RPE results for the trans. (mm).
Sequence ID Rel-Pose Raw Abs-Pose Refined Abs-Pose PGO Pose
RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE. RPE.
trans.max trans.mean trans.median trans.max trans.mean trans.median trans.max trans.mean trans.median trans.max trans.mean trans.median
Seq0 70.27 12.72 9.16 18.49 3.80 3.25 18.45 3.72 3.17 13.98 3.73 3.18
Seq1 56.04 10.16 7.97 18.60 2.17 1.66 18.39 2.11 1.64 17.18 2.12 1.64
Seq2 656.47 16.70 9.84 11.71 2.89 2.38 11.67 2.87 2.33 11.67 2.88 2.33
Seq3 37.90 8.66 7.70 18.67 3.31 2.10 18.74 3.28 2.05 18.74 3.28 2.05
Seq4 32.24 9.92 8.79 13.85 2.68 2.13 13.83 2.65 2.07 12.63 2.65 2.07
Seq5 41.07 6.98 6.09 6.49 1.28 1.03 6.51 1.27 1.00 6.51 1.27 1.00
Seq6 34.65 8.84 7.37 15.16 3.50 3.14 15.12 3.47 3.06 15.12 3.49 3.11
Seq7 35.43 8.35 6.72 18.15 2.37 1.51 18.16 2.27 1.46 17.85 2.27 1.46
Seq8 95.26 7.78 6.16 11.75 1.53 1.14 11.76 1.49 1.13 11.76 1.49 1.13
Seq9 27.24 6.78 6.05 10.58 1.96 1.72 10.23 1.94 1.69 10.23 1.94 1.69
Avg. 108.66 9.69 7.59 14.35 2.55 2.01 13.35 2.51 1.96 12.63 2.51 1.97
(a) Seq0 - Mocap
(b) Seq1 - Mocap
(c) Seq2 - Mocap
(d) Seq3 - Mocap
(e) Seq4 - Mocap
(f) Seq0 - PGO
(g) Seq1 - PGO
(h) Seq2 - PGO
(i) Seq3 - PGO
(j) Seq4 - PGO
Figure 5: Comparison of algorithm performance across Seq0 - Seq4 sequences. The top row shows Mocap results, and the bottom row shows PGO results for the same sequences.
(a) Seq5 - Mocap
(b) Seq6 - Mocap
(c) Seq7 - Mocap
(d) Seq8 - Mocap
(e) Seq9 - Mocap
(f) Seq5 - PGO
(g) Seq6 - PGO
(h) Seq7 - PGO
(i) Seq8 - PGO
(j) Seq9 - PGO
Figure 6: Comparison of algorithm performance across Seqs5 - Seqs9 sequences. Top row shows Mocap results, and the bottom row shows PGO results for the same sequences.
VI-A3 Results
Here are our experimental comparison results, where we report the mean, max, and median for ATE shown in table VII and RPE shown in table VIII and table IX. Additionally, we randomly select several frames to present qualitative visualization results. The results show that our final ATE and RPE are very close to the ground truth, with an average translation error of less than $3$ mm and an average rotation error of less than $0.2$ degree, indicating quite good quality. Furthermore, while the overall results after pose graph optimization didn’t change significantly, the maximum error decreases. This aligns with the design goal of this module: effectively compensating for poor estimations in Refined Abs-Pose and preserving the good ones. From the perspective of relative pose estimation itself, its drift is still noticeable, and its overall accuracy cannot match that of Refined Abs-Pose. Therefore, it can only serve to compensate for low-quality absolute poses.
VI-B Analysis of Unseen Object Pose Estimation Methods
Figure 7: FoundPose[ornek2024foundpose]* failure case.
Illustrated in fig. 7, [ornek2024foundpose]* heavily relies on features extracted by DINOv2 [oquab2023dinov2] to establish 2D-3D correspondences, which makes it susceptible to pose estimation errors for low-texture or highly symmetric objects. fig. 7 shows that the object (a frying pan) is textureless, and during the detection process, the right-side handle was mistakenly ignored. As a result, the object was erroneously treated as symmetric, leading to incorrect pose estimates.
Figure 8: FoundationPose[wen2024foundationpose] failure case.
As shown in fig. 8, we directly employ the model-based [wen2024foundationpose]. It primarily relies on the input RGB-D data. When the camera undergoes rapid motion, causing motion blur, the pose estimation results becomes error-prone as well. This issue is especially prominent for flat, textureless, and symmetric objects, where the network may incorrectly estimate the orientation of the coordinate axes.
Figure 9: GigaPose [nguyen2024gigapose] failure case.
The performance of GigaPose [nguyen2024gigapose] benefits from its time-intensive pre-processing step, where it renders the input CAD model extensively before performing pose estimation. This process allows it to establish better correspondences during object pose estimation. However, since the method relies solely on RGB information and performs feature matching across different viewpoints of the CAD model, it faces significant challenges when dealing with occlusions or textureless objects as shown in fig. 9, such limitations result in a failure case where the person’s hand holding the mouse causes the estimation to break down.
(a) MegaPose (RGB)
(b) MegaPose (RGB-D)
Figure 10: MegaPose[labbe2022megapose] failure case.
As shown in fig. 10, here we evaluate both the RGB-based and RGB-D-based variants of [labbe2022megapose]. This method takes a cropped image (RGB or RGB-D) as input along with the corresponding CAD model. Multiple poses are rendered from the CAD model and then passed to the coarse estimator to obtain the initial poses, which are subsequently refined by the refinement network to produce the final outputs. Note that the final results heavily depend on the coarse estimator network. If the initial pose estimation is inaccurate, then the refinement stage is generally unable to recover the correct poses. Therefore, both RGB and RGB-D approaches are highly dependent on the quality of the multiview input poses. For geometrically symmetric or regular-shaped objects such as the juice box shown in fig. 10, the use of multiple poses from the CAD model can easily lead to incorrect initial guesses. Consequently, failure cases are observed in both the RGB-based and RGB-D-based versions of the method, as illustrated in the figure fig. 10.
Figure 11: SAM6D[lin2024sam] failure case.
SAM6D [lin2024sam] is a multi-stage method. It first segments any objects using SAM[kirillov2023segment], and then associate the segmentation mask to the target object. And through the [qin2023geotransformer] to estimate the object pose. Incorrect registration may happen in symmetric and textureless object cases, resulting in failure cases such as the one shown in fig. 11.
VI-C Analysis of Object Pose Tracking Methods
Figure 12: MaskFusion [runz2018maskfusion] failure case.
For [runz2018maskfusion], to ensure a fair benchmark, we replaced its original mask results with masks generated by the state-of-the-art method SAM2 [ravi2024sam]. However, when intensive motion occurs, the object tracking algoritm tends to reinitialize its state. Consequently, this leads to noticeable drift. An example failure case is shown in fig. 12
Figure 13: BundleTrack [wen2021bundletrack] failure case.
[wen2021bundletrack] takes as input an RGB-D sequence along with object segmentation results. It then uses traditional keypoint matching methods to estimate the relative poses between consecutive frames, which are subsequently optimized through an object pose graph. However, as fig. 13 shows, its main drawback is that for objects with sparse texture, the relative pose estimations often become inaccurate, leading to accumulated drift in subsequent frames.
Figure 14: BundleSDF [wen2023bundlesdf] failure case.
fig. 14) demonstrates overall robust performance without obvious drift or failure cases. However, it still faces considerable challenges when dealing with textureless and symmetric objects. Moreover, as [wen2023bundlesdf] is a Neural Object Field-based approach, it simultaneously estimates the object’s 6DoF pose and reconstructs the object’s mesh. Consequently, the runtime for processing an entire sequence is considerably limited.
Figure 15: DROID-SLAM [teed2021droid] failure case.
[teed2021droid], as a popular SLAM framework, takes RGB-D inputs and was originally proposed for camera pose tracking and scene reconstruction. When provided with segmented images and fed into its DBA layer, it operates in an object-centric setting, allowing us to obtain the object pose for each frame. However, since the masked objects provide limited input information, any error in camera pose estimation can lead to significant drift(Fig.15).
Figure 16: FoundationPose [wen2024foundationpose] (tracking) failure case.
[wen2024foundationpose] supports both object pose estimation and object pose tracking. We benchmark its tracking mode. As shown in fig. 16, the tracking results in drift failure and tracking loss in rapid motion.
VI-D Analysis of Category-level Object Estimation Methods
(a) Original AG-Pose [lin2024instance].
(b) Fine-tuned AG-Pose [lin2024instance].
Figure 17: AG-Pose [lin2024instance].
(a) Original SecondPose [chen2024secondpose]
(b) Fine-tuned SecondPose [chen2024secondpose].
Figure 18: SecondPose [chen2024secondpose].
VI-D1 Fine-tuning
Moreover, under our Dynamic Object with Moving Camera scenarios, the dataset presents different challenges compared to static real-world datasets such as REAL275 [wang2019normalized] or HouseCat6D [jung2024housecat6d]. Our dataset captures novel viewpoints induced by object rotations along different axes while being held in hand—viewpoints that are difficult to observe in static scenes. These challenges manifest in pose estimation failures for various methods, such as GCASP [li2023generative] when a bowl is lifted upward, AG-Pose [lin2024instance] and GenPose [zhang2023generative] when the camera is moved upward, and DiffusionNOCS [ikeda2024diffusionnocs] when a can is lifted vertically. Furthermore, in our fine-tuning experiments (see Section  III), it is evident that fine-tuning on our dataset leads to substantial improvements. For example, in Fig.17 and Fig.18, when comparing SecondPose [chen2024secondpose] and AG-Pose [lin2024instance] before and after fine-tuning, the previously mentioned pose estimation failures caused by vertical lifting of objects are significantly mitigated.
(a) NOCS [wang2019densefusion].
(b) NOCS (MV-ROPE) [yang2024mv].
(c) DiffusionNOCS [ikeda2024diffusionnocs].
(d) AG-Pose [lin2024instance].
(e) SecondPose [chen2024secondpose].
(f) IST-Net [liu2023net]
(g) VI-Net [lin2023vi].
(h) GCASP [li2023generative].
(i) GenPose [zhang2023generative].
Figure 19: COPE methods failure cases.
VI-E COPE methods failure cases.
[wang2019normalized] predicts normalized object points and then apply the Umeyama algorithm [umeyama1991least] to recover the object pose and size. However [umeyama1991least] requires high-quality NOCS map and depth map to recover accurate object pose and size, otherwise, its accuracy is limited. Note that in the original NOCS implementation, its object detection and NOCS map prediction module is coupled. For fairness in benchmarking, we adopt the settings from MV-ROPE [yang2024mv], where consistent segmentation results similar to other benchmarked methods are provided as input.
Overall, the methods that rely on predicting a NOCS map, such as [wang2019normalized, ikeda2024diffusionnocs, yang2024mv] tend to exhibit significantly higher translation errors compared to methods that directly predict translation, such as [liu2023net, li2023generative, lin2023vi, chen2024secondpose, lin2024instance, zhang2023generative]. This performance gap can be attributed to the fact that most of the latter methods, including [liu2023net, lin2023vi, chen2024secondpose, lin2024instance], perform translation prediction using PointNet-based architectures[qi2017pointnet] and decouple it from rotation prediction. Compared to rotation estimation, translation prediction techniques are relatively more mature and stable.
Regarding COPE algorithms, we found that challenging cases tend to concentrate in two categories: the camera category and the mug category, so we focused our analysis there.
  1. 1.
    The camera category has long been a known challenge in COPE tasks due to the large intra-class variation among commonly seen camera models. Cameras in real-world scenarios exhibit a wide range of appearances, posing significant challenges for the generalization ability of pose estimation networks.
  2. 2.
    The mug category, by contrast, tends to have more consistent geometric shapes. However, failure cases often occur when a mug is being held by a human hand. Such occlusion frequently leads to errors in rotation estimation. Another possible cause of failure is when the mug’s opening is not visible in the input image, under such conditions, it becomes difficult for the network to distinguish the upright orientation of the mug.
Figure 20: Viewpoint coverage map. Most of our benchmarked COPE algorithm is trained on NOCS-6D [wang2019densefusion] dataset. Differ from NOCS-6D [wang2019normalized], our viewpoint coverage is more balanced between top-down and bottom-up views. Note that the left figure is referred to [jung2024housecat6d].
Another issue is that most benchmarked COPE methods are trained in NOCS [wang2019normalized] dataset, whose pose distribution is different from ours, which can be another reason for the failure cases in our test split.
[…]
[heading=bibintoc,title=References]